Show from national income accounting that
a. An increase in taxes (while transfers remain constant) must imply a change in net
exports, government purchases, or the saving-investment balance.
b. An increase in disposable personal income must imply an increase in consumption or an
increase in saving.
c. An increase in both consumption and saving must imply an increase in disposable
income.
[ For both parts b and c assume there are no interest payments by households or transfer pay￾ments to foreigners.]
1
Expert's answer
2021-08-13T10:23:14-0400

а. We calculate taxes based on the economy's identification with the state and international commerce.

S - I = (G + TR - TA) + NX

TA = G + TR + NX + I - S

If taxes go up, the other side of the equation should alter as well.

As a result, the change is a constant transition: a higher G, higher NX, higher I, or lower S.

b. Based on the equation YD = S + C criteria, an increase in disposable income might be saved or consumed.

c. An rise in consumption and savings should increase disposable income, as predicted by the equation above.